# On structural contraction of biological interaction networks

### v3

@@ 7 Conclusions
where $\Lambda_{11},\Lambda_{12},\Lambda_{21},\Lambda_{22}\in\mathbb{R}^{n\times n}$. By the statement in Theorem 2.14, $\Lambda_{11},\Lambda_{22}$ are Metzler and $\Lambda_{12},\Lambda_{21}$ are non-nonnegative. In addition, (16) can be written as:
@@
Note that we can assume, w.l.o.g, that $\Lambda_{22}=\Lambda_{11}$, $\Lambda_{21}=\Lambda_{12}$. Hence, we define $\Lambda_{\ell}:=\Lambda_{11}-\Lambda_{12}$. It remains to show that $\mu_{\infty}(\Lambda_{\ell})\leq 0$. To that end, recall that $\mu_{\infty}(\Lambda_{\ell})=\max_{i}\lambda_{ii}^{(\ell)}+\sum_{j\neq i}|\lambda_{ij}^{(\ell)}|$. Hence, we write:
@@
where the inequality ()($\star$)( ⋆ ) follows since $\lambda_{ii}^{12},\lambda_{ij}^{11},\lambda_{ij}^{12}$ are nonnegative, and equality ()($\clubsuit$)( ♣ ) follows from $\tilde{\Lambda}_{\ell}\mathbf{1}=0$. Hence, we have found a matrix $\Lambda_{\ell}$ that satisfy $\mu_{\infty}(\Lambda_{\ell})\leq 0$.
Next, we show the other direction. Fix $\ell$. Assume we have a matrix $\Lambda_{\ell}$ that satisfies $\mu_{\infty}(\Lambda_{\ell})\leq 0$. Hence, for all $i$ we have $\lambda_{ii}^{(\ell)}\leq 0$ and $\sigma_{i}:=-(\lambda_{ii}^{(\ell)}+\sum_{j}|\lambda_{ij}^{(\ell)}|)\geq 0$. We are going to construct two matrices $\Lambda_{11},\Lambda_{12}$ elementwise as follows assuming $i\neq j$: $\lambda_{ij}^{(11)}:=\max\{\lambda_{ij}^{(\ell)},0\}+\sigma_{i}/{(2n)},\lambda_{ij}^{(12)}:=\max\{-\lambda_{ij}^{(\ell)},0\}+\sigma_{i}/{(2n)}$, $\lambda_{ii}^{(12)}:=\sigma_{i}/{(2n)}$, $\lambda_{ii}^{(11)}:=\lambda_{ii}^{(\ell)}+\sigma_{i}/{(2n)}$, where the latter is non-positive since $\lambda_{ii}^{(\ell)}+\sigma_{i}/{(2n)}\leq\lambda_{ii}^{(\ell)}+\sigma_{i}\leq\lambda_{ii}^{(\ell)}-\lambda_{ii}^{(\ell)}=0$. Finally, we define $\Lambda_{22}:=\Lambda_{11},\Lambda_{21}:=\Lambda_{12}$. Define $\tilde{\Lambda}_{\ell}$ as in (33) which can be verified to be a Metzler matrix satisfying $\tilde{\Lambda}_{\ell}\mathbf{1}=0$ and (34)-(35). $\square$
@@
Remember that $\rho_{1},..,\rho_{s}$ correspond to the nonzero entries of $\partial R/\partial x$ which are strictly positive in $\mathbb{R}_{+}^{n}$ by AK3, hence the proof can be accomplished by showing that $\mu_{\infty}\left(\sum_{\ell=1}^{s}\rho_{\ell}\Lambda_{\ell}\right)<0$ for any positive $\rho_{1},..,\rho_{s}$. Assume $\Theta\in\mathbb{R}^{m\times m}$. Since $\Theta$ might have zero rows, the function can be written (by reordering species if necessary) as ${\widetilde{V}}(r)=\|[\Theta_{d},O]\Gamma r\|_{\infty}$, where $\Theta_{d}$ is an $n_{d}\times n_{d}$ positive diagonal matrix for some $n_{d}\leq n$. Hence, ${\widetilde{V}}(r)=\max_{i\in\{1,..,n_{d}\}}\gamma_{i}^{T}r$. To show the statement, it is sufficient to show that for each $i\in\{1,..,n_{d}\}$, there exists $\ell\in\{1,..s\}$ such that $\lambda_{\ell}^{(ii)}+\sum_{j\neq i}|\lambda_{\ell}^{(ik)}|<0$.
By AS1, $\gamma_{i}$ is orthogonal to a positive vector, and hence must have a strictly negative element, say $\gamma_{ij}$. Hence, this implies that $X_{i}$ is a reactant of $\mathbf{R}_{j}$, and $\partial R_{j}/\partial x_{i}$ is nonzero. Let $\Lambda_{\ell}$ the matrix (defined in Corollary 8) corresponding to the reactant-reactant pair $(X_{i},{\rm\bf R}_{j})$. Therefore, it can be seen it is possible to choose the $i$th row of $\Lambda_{\ell}$ as follows: $\lambda_{\ell}^{(ii)}=-\theta_{ii}$ and $\lambda_{\ell}^{(ij)}=0,i\neq j$. Hence, $\lambda_{\ell}^{(ii)}+\sum_{j\neq i}|\lambda_{\ell}^{(ij)}|=-\theta_{ii}<0$. $\square$
@@ Acknowledgments
M.A.A and E.D.S. partially supported by grants AFOSR FA9550-21-1-0289 and NSF/DMS-2052455.

### v4

@@ 7 Conclusions
where $\Lambda_{11},\Lambda_{12},\Lambda_{21},\Lambda_{22}\in\mathbb{R}^{{m\times m}}$. By the statement in Theorem 2.14, $\Lambda_{11},\Lambda_{22}$ are Metzler and $\Lambda_{12},\Lambda_{21}$ are non-nonnegative. In addition, (16) can be written as:
@@
Note that we can assume, w.l.o.g., that $\Lambda_{22}=\Lambda_{11}$, $\Lambda_{21}=\Lambda_{12}$. Hence, we define $\Lambda_{\ell}:=\Lambda_{11}-\Lambda_{12}$. It remains to show that $\mu_{\infty}(\Lambda_{\ell})\leq 0$. To that end, recall that $\mu_{\infty}(\Lambda_{\ell})=\max_{i}\lambda_{ii}^{(\ell)}+\sum_{j\neq i}|\lambda_{ij}^{(\ell)}|$. Hence, we write:
@@
where the inequality $(\star)$ follows since $\lambda_{ii}^{12},\lambda_{ij}^{11},\lambda_{ij}^{12}$ are nonnegative, and equality $(\clubsuit)$ follows from $\tilde{\Lambda}_{\ell}\mathbf{1}=0$. Hence, we have found a matrix $\Lambda_{\ell}$ that satisfies $\mu_{\infty}(\Lambda_{\ell})\leq 0$.
Next, we show the other direction. Fix $\ell$. Assume we have a matrix $\Lambda_{\ell}$ that satisfies $\mu_{\infty}(\Lambda_{\ell})\leq 0$. Hence, for all $i$ we have $\lambda_{ii}^{(\ell)}\leq 0$ and ${\eta}_{i}:=-(\lambda_{ii}^{(\ell)}+\sum_{j}|\lambda_{ij}^{(\ell)}|)\geq 0$. We are going to construct two matrices $\Lambda_{11},\Lambda_{12}$ elementwise as follows assuming $i\neq j$: $\lambda_{ij}^{(11)}:=\max\{\lambda_{ij}^{(\ell)},0\}+{\eta}_{i}/{(2n)},\lambda_{ij}^{(12)}:=\max\{-\lambda_{ij}^{(\ell)},0\}+{\eta}_{i}/{(2n)}$, $\lambda_{ii}^{(12)}:={\eta}_{i}/{(2n)}$, $\lambda_{ii}^{(11)}:=\lambda_{ii}^{(\ell)}+{\eta}_{i}/{(2n)}$, where the latter is non-positive since $\lambda_{ii}^{(\ell)}+{\eta}_{i}/{(2n)}\leq\lambda_{ii}^{(\ell)}+{\eta}_{i}\leq\lambda_{ii}^{(\ell)}-\lambda_{ii}^{(\ell)}=0$. Finally, we define $\Lambda_{22}:=\Lambda_{11},\Lambda_{21}:=\Lambda_{12}$. Define $\tilde{\Lambda}_{\ell}$ as in (33) which can be verified to be a Metzler matrix satisfying $\tilde{\Lambda}_{\ell}\mathbf{1}=0$ and (34)-(35). $\blacksquare$
@@
Remember that $\rho_{1},\dots,\rho_{s}$ correspond to the nonzero entries of $\partial R/\partial x$ which are strictly positive in $\mathbb{R}_{+}^{n}$ by AK3, hence the proof can be accomplished by showing that $\mu_{\infty}\left(\sum_{\ell=1}^{s}\rho_{\ell}\Lambda_{\ell}\right)<0$ for any positive $\rho_{1},\dots,\rho_{s}$. Assume $\Theta\in\mathbb{R}^{m\times m}$. Since $\Theta$ might have zero rows, the function can be written (by reordering species if necessary) as ${\widetilde{V}}(r)=\|[\Theta_{d},O]\Gamma r\|_{\infty}$, where $\Theta_{d}$ is an $n_{d}\times n_{d}$ positive diagonal matrix for some $n_{d}\leq n$. Hence, ${\widetilde{V}}(r)=\max_{i\in\{1,\dots,n_{d}\}}\gamma_{i}^{T}r$. To show the statement, it is sufficient to show that for each $i\in\{1,\dots,n_{d}\}$, there exists $\ell\in\{1,\dots,s\}$ such that $\lambda_{\ell}^{(ii)}+\sum_{j\neq i}|\lambda_{\ell}^{(ik)}|<0$.
By AS1, $\gamma_{i}$ is orthogonal to a positive vector, and hence must have a strictly negative element, say $\gamma_{ij}$. Hence, this implies that $X_{i}$ is a reactant of $\mathbf{R}_{j}$, and $\partial R_{j}/\partial x_{i}$ is nonzero. Let $\Lambda_{\ell}$ be the matrix (defined in Corollary 8) corresponding to the reactant-reaction pair $(X_{i},{\rm\bf R}_{j})$. Therefore, it can be seen it is possible to choose the $i$th row of $\Lambda_{\ell}$ as follows: $\lambda_{\ell}^{(ii)}=-\theta_{ii}$ and $\lambda_{\ell}^{(ij)}=0,i\neq j$. Hence, $\lambda_{\ell}^{(ii)}+\sum_{j\neq i}|\lambda_{\ell}^{(ij)}|=-\theta_{ii}<0$. $\blacksquare$
@@ Acknowledgments
M.A.A and E.D.S. acknowledge the support of grants AFOSR FA9550-21-1-0289 and FA9550-22-1-0316, and NSF/DMS-2052455.